OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. 1, -1, -3, -5, ...

OpenStudy (anonymous):

t(n)=1+(n-1)(-2) t(n)=3-2n

OpenStudy (anonymous):

here a site to help www.mathway.com

OpenStudy (anonymous):

thankss